The average repair cost of a random sample of 100 refrigerators is $150 with a standard deviation $15. If (a,b) is the 80% confidence interval for the population mean cost constructed by using this sample, then

a=[a] enter a dollar amount, round to the nearest cents.

b=[b] enter a dollar amount, round to the nearest cents.
Answer #1

n-100 TO S given T =$150 s = $15 construct 80%. c. I. for mean. It * T tt at (n-1)=(100-1) =gg degree of Freedom +* = 1.2902